A heating element is a (nichrome) wire with cross-sectional area of 2 x 10^-7 m^2 and is 1.3 m long.  The material has resistivity of 4 x 10^-5 ohm-m at 200°C and a temperature coefficient of 3 x 10^-2 /°C.  Find the resistance of the element at 350°C


1
Expert's answer
2021-04-13T06:28:33-0400

The resistance is


"R=\frac{\rho_{350} L}{A}=\frac{\rho_{200}(1+\alpha \Delta T) L}{A},\\\space\\
R=\frac{4·10^{-5}(1+0.03(350-200)) 1.3}{2·10^{-7}}=1430\space\Omega."